5.8 A car is driving along a highway that goes over a hill. The top of the hill has a radius of curvature of 45 m. What is te maximum safe speed for going over the hilltop?
Answer #1

Fg = m*v^2/R

m*g = mv^2/R


V = sqrt(Rg) = 21 m/s